Knowing that for the extruded beam shown the allowable stress is 120 MPa in tension and 147 MPa in compression, determine the largest couple M that can be applied.

The largest couple M that can be applied is 13.83 13.83 Incorrect kN·m.
Answer #1

48 mmo I 96mm Monment of Centroide from top A = 48x96. i 9 = 96/2=48 Az = 36x36, 42 = 48+ 36/2 = 66 mm x ge Ayi-A292 48x96x48